# Searching for emergent spacetime in spin glasses An algebraic perspective

### v1

@@ 3.2 A bulk from the spectral function
It is evident that the depth parameter $\mathcal{T}$} of Def. 1 is given by the length of the smallest time interval, so that the algebra of the timeband has no relative commutant anymore. The relevance of the spectral functions for emergent bulk properties was already conjectured in [48, 49]. In the bulk language, this corresponds to a bulk reconstruction of a full Cauchy surface. This is due to the fact that if one can reconstruct a full Cauchy surface, no nontrivial operator can commute with all operators in the algebra anymore, since the algebra associated to the Cauchy surface contains all operators [^7] This hinges technically on the statement that the timeband algebras are factors, which means that every operator that is not the identity and commutes with every operator in the algebra is not itself part of the algebra.. GL then proved the following statements (which we repeat in an informal fashion).
@@ 4.3 The SU($M$) Heisenberg model
Increasing $\kappa$ even more moves us into the spin glass phase. For fixed $J=0.2$ this part of the spin glass phase corresponds to the semiclassical regime [34]. In this regime, we find that the support of $\rho(\omega)$ is compact by the emergence of a kink, thus neglecting the possibility of a semiclassical emergent radial direction from Theorem 1. The behavior is quite similar to the one observed in the spin glass phase of the spherical $p$-spin model, with the difference that there is a part of exponential decay for small frequencies, that leads to a kink and a sudden, almost vertical, decay. This is followed by a superexponential tail that becomes exponential. Due to the order of magnitude drop at low frequencies, that does not analogously appear in the spin liquid phase, we consider it as a sign of compact support. In addition, we see that a gap opens up for increasing $\kappa$. It would be interesting whether the limit $\kappa\rightarrow\infty$ exists, since the support appears to increase with increasing $\kappa$.
@@ 5 No detectable emergence for low-energy observables
Proof: Assume that a function $g(t)$ exists, so that the smeared operator $O(g)$ is in the relative commutant for the timeband algebra associated to $\Delta T$. Following the discussion below (38), the relevant function to consider is the Fourier transform of $g(\omega)\rho(\omega)$, which has to vanish inside $\Delta T$, i.e.
${\mathcal{F}}(t)=\int d\omega\rho(\omega)g(\omega)e^{i\omega t}.$ (82)
Since $\rho(\omega)$ is exponentially decaying, multiplying it with any polynomial function will still lead to a function that is in $L^{1}(\mathbb{R})$. In particular, ${\mathcal{F}}(t)$ defines a function that is analytic in the strip $\mathbb{R}+i[-\alpha,\alpha]$. If ${\mathcal{F}}(t)$ vanishes inside $\Delta T$, by analyticity it has to be zero everywhere. Therefore no such $g(t)$ exists. ∎
@@ Acknowledgments
We thank Elliott Gesteau, Zhuo-Yu Xian, Eric Anschuetz and Alexander Jahn for fruitful discussions and comments on the draft. We thank Ben Karlsberg for his collaboration during the initial stages of this project. We thank Felix Haehl for communication on their implementation of the numerical procedure. This work has been supported by the Einstein Research Unit on Quantum Devices and Berlin Quantum.
@@ B.5 The saddle-point approximation
The next step is to eliminate the matrices $Q_{rr^{\prime}}(t-t^{\prime})$ and $\Sigma_{rr^{\prime}}(\tau,\tau^{\prime})$ from the replica effective action 137 above, via the saddle-point approximation, which is exact for $N\rightarrow\infty$, as we also mentioned above. First, we vary the effective action 137 over $Q_{rr^{\prime}}(\tau-\tau^{\prime})$ and obtain
@@
Next, we vary the effective action 137 over $\Sigma_{rr^{\prime}}(\tau,\tau^{\prime})$. Let us define $K_{rr^{\prime}}(\tau,\tau^{\prime})=\delta_{rr^{\prime}}\delta(\tau-\tau^{\prime})(\partial_{\tau^{\prime}}+i\lambda_{r}(\tau))$}
@@ B.6 Schwinger-Dyson equations
We can derive the Schwinger-Dyson equations by varying the effective action 148 with respect to the bosonic fields, $G_{rr^{\prime}}$:
@@ B.7 The replica symmetry breaking ansatz (RSB-1)
In [97], the authors argued that the matrix elements $Q_{rr^{\prime}}(\tau-\tau^{\prime})$ in 130 are time-dependent only for $r=r^{\prime}$, meaning that all off-diagonal replica terms are time-independent. For the purposes of our analysis, we will make the same assumption for the elements of $G_{rr^{\prime}}(\tau-\tau^{\prime})$. In particular we will assume that the SG solutions of the models we study, both the $p$-spin and SU(M)(M) Heisenberg models, satisfy the replica symmetry breaking ansatz at one-step (RSB-1). This means that there is an additional block-diagonal structure that the $G_{rr^{\prime}}(\tau-\tau^{\prime})$ matrices should satisfy, as we show in (152). The RSB-1 has been already used for the $p$-spin model in [89, 36, 30], as well as for the SU(M)(M) Heisenberg model in [34]. An example of such an matrix ansatz for block-diagonal matrices of dimension $3\times 3$ is:
@@
Next, we insert the ansatz above into the effective action 148, assuming that also $\lambda_{r}(\tau)$ is independent of the replica index:
@@
We can now compute the equation of motion (EoM) for $s$, the parameter that determines the structure of the ansatz outside the block-matrices. Once again, we need to vary 154 over the corresponding parameter. Then, we have:
@@ B.8 Equations of motion in momentum space
Under this convention, the time-dependent diagonal elements of $\mathbf{G}(\tau-\tau^{\prime})$ in 152, i.e., the $g(\tau,\tau^{\prime})$ terms, can be transformed as:
@@
Plugging 158 in the effective action 156, we can finally get its form in momentum space, as:
@@
Every point on the phase diagram of the SU(M)(M) Heisenberg model satisfies all EoMs 171, 164, 165 and 166 simultaneously. For the spin liquid phase we expect $u=0$ and $m=1$ for all $J$ and $\kappa$. On the other hand, the spin glass phase is characterized by $0<u\leq\kappa$ and $0<m<1$. The bound for $u$ arises from the constraint Eq. (76) and for integer $n$, $m$ is also an integer that describes the size of the diagonal blocks in the ansatz. In the limit $n\rightarrow 0$, $m$ becomes a continuous number between 0 and $1$ [30, 36]. In order to be able to find the correct solutions $\hat{g}(k)$ numerically for any point on the phase diagram, we need first to massage the EoMs into a more convenient form.
@@ B.9 Lorentzian two point function
The relation above is quite useful, as it will let us compute $g^{R}(\omega)$, from the values $\{u,m,g(k=0)\}$ that satisfy the triplet of Euclidean equations of motion 164, 165 and 166, derived above. In particular, the analytic continuation of (171) is the following:
@@ Appendix C Details on the spectral function
where the $\pm$-sign in the last line defines the fermionic and bosonic Boltzmann factors.
@@ Appendix D Analytic continuation of equations of motion
As we saw in 194, the retarded Green’s function is related to the euclidean Green’s function by an analytic continuation of the form $\omega\rightarrow-i\omega$. We can apply the same procedure to the Schwinger-Dyson equations to obtain the real time equations. which in turn allows us to find an expression for the self energy $\Sigma(\tau)$

### v2

@@ 3.2 A bulk from the spectral function
One can show [35], that for generalized free fields, the depth parameter $\mathcal{T}$} of Def. 1 is given by the length of the smallest time interval, so that the algebra of the timeband has no relative commutant anymore. The relevance of the spectral functions for emergent bulk properties was already conjectured in [28, 29]. In the bulk language, this corresponds to a bulk reconstruction of a full Cauchy surface. This is due to the fact that if one can reconstruct a full Cauchy surface, no nontrivial operator can commute with all operators in the algebra anymore, since the algebra associated to the Cauchy surface contains all operators [^7] This hinges technically on the statement that the timeband algebras are factors, which means that every operator that is not the identity and commutes with every operator in the algebra is not itself part of the algebra.. GL then proved the following statements (which we repeat in an informal fashion).
@@ 4.3 The SU($M$) Heisenberg model
Increasing $\kappa$ even more moves us into the spin glass phase. For fixed $J=0.2$, this part of the spin glass phase corresponds to the semiclassical regime [34]. In this regime, for finite $\kappa$, we observe the emergence of a kink, which can be seen as an almost vertical decay of the spectral function in the logarithmic plot 10 d). After this drop, $\rho(\omega)$ exhibits a superexponential decay that becomes an exponential tail. The existence of a kink is quite similar to the one observed in the spin glass phase of the spherical $p$-spin model, with the difference that the function exhibits a regime of exponential decay that precedes the kink. Due to the order of magnitude drop, that does not analogously appear in the spin liquid phase, we consider it as a sign of compact support at finite $\kappa$. In addition, we note that as $\kappa$ increases, the spectral function tends to zero at low frequencies, which we interpret as the emergence of a spectral gap (see Fig. 10 c) for $\kappa=8$). Furthermore, with increasing $\kappa$ the support appears to become larger, which is indicated by the position of the kink moving to larger frequencies. It would be interesting to study, whether the limit $\kappa\rightarrow\infty$ exists, since by extrapolation of the finite $\kappa$ data, this limit might exhibit noncompact support.
@@ 5 No detectable emergence for low-energy observables
Proof.
Assume that a function $g(t)$ exists, so that the smeared operator $O(g)$ is in the relative commutant for the timeband algebra associated to the interval $\Delta T$ and its Fourier transform $g(\omega)={\mathcal{O}}(\omega^{k})$ for some constant $k>0$. Following the discussion below Eq. (38), the relevant function to consider is the Fourier transform of $g(\omega)\rho(\omega)$, which has to vanish inside $\Delta T$ for $O(g)$ to be in the relative commutant, i.e.
${\mathcal{F}}(t)=\int d\omega\rho(\omega)g(\omega)e^{i\omega t}=0,\ \forall t\in\Delta T.$ (82)
Since $\rho(\omega)={\mathcal{O}}(\exp(-\alpha\omega^{\beta}))$, the product $g(\omega)\rho(\omega)$ with polynomially growing $g(\omega)$ is in $L^{1}(\mathbb{R})$ and in particular we have
$g(\omega)\rho(w)={\mathcal{O}}(\omega^{k}\exp(-\alpha\omega^{\beta})).$ (83)
Because of this decay, the Fourier transform ${\mathcal{F}}(t)$ defines a function that is analytic in the interior of the strip $\mathbb{R}+i[-\alpha,\alpha]$ for $\beta=1$ and an entire function for $\beta>1$. Therefore, if ${\mathcal{F}}(t)$ vanishes inside $\Delta T$, by analyticity it has to be zero everywhere, which implies that $g(\omega)=0$ everywhere. Therefore no such nonzero $g(t)$ exists. ∎
@@ Acknowledgments
We thank Elliott Gesteau, Zhuo-Yu Xian, Eric Anschuetz and Alexander Jahn for fruitful discussions and comments on the draft. We thank Ben Karlsberg for his collaboration during the initial stages of this project. We thank Felix Haehl for communication on their implementation of the numerical procedure. We gratefully acknowledge computing time on the HPC clusters at the physics department of FU Berlin. This work has been supported by the Einstein Research Unit on Quantum Devices and Berlin Quantum.
@@ B.5 The saddle-point approximation
The next step is to eliminate the matrices $Q_{rr^{\prime}}(t-t^{\prime})$ and $\Sigma_{rr^{\prime}}(\tau,\tau^{\prime})$ from the replica effective action 138 above, via the saddle-point approximation, which is exact for $N\rightarrow\infty$, as we also mentioned above. First, we vary the effective action 138 over $Q_{rr^{\prime}}(\tau-\tau^{\prime})$ and obtain
@@
Next, we vary the effective action 138 over $\Sigma_{rr^{\prime}}(\tau,\tau^{\prime})$. Let us define $K_{rr^{\prime}}(\tau,\tau^{\prime})=\delta_{rr^{\prime}}\delta(\tau-\tau^{\prime})(\partial_{\tau^{\prime}}+i\lambda_{r}(\tau))$}
@@ B.6 Schwinger-Dyson equations
We can derive the Schwinger-Dyson equations by varying the effective action 149 with respect to the bosonic fields, $G_{rr^{\prime}}$:
@@ B.7 The replica symmetry breaking ansatz (RSB-1)
In [9], the authors argued that the matrix elements $Q_{rr^{\prime}}(\tau-\tau^{\prime})$ in 131 are time-dependent only for $r=r^{\prime}$, meaning that all off-diagonal replica terms are time-independent. For the purposes of our analysis, we will make the same assumption for the elements of $G_{rr^{\prime}}(\tau-\tau^{\prime})$. In particular we will assume that the SG solutions of the models we study, both the $p$-spin and SU(M)(M) Heisenberg models, satisfy the replica symmetry breaking ansatz at one-step (RSB-1). This means that there is an additional block-diagonal structure that the $G_{rr^{\prime}}(\tau-\tau^{\prime})$ matrices should satisfy, as we show in (153). The RSB-1 has been already used for the $p$-spin model in [17, 19, 5], as well as for the SU(M)(M) Heisenberg model in [34]. An example of such an matrix ansatz for block-diagonal matrices of dimension $3\times 3$ is:
@@
Next, we insert the ansatz above into the effective action 149, assuming that also $\lambda_{r}(\tau)$ is independent of the replica index:
@@
We can now compute the equation of motion (EoM) for $s$, the parameter that determines the structure of the ansatz outside the block-matrices. Once again, we need to vary 155 over the corresponding parameter. Then, we have:
@@ B.8 Equations of motion in momentum space
Under this convention, the time-dependent diagonal elements of $\mathbf{G}(\tau-\tau^{\prime})$ in 153, i.e., the $g(\tau,\tau^{\prime})$ terms, can be transformed as:
@@
Plugging 159 in the effective action 157, we can finally get its form in momentum space, as:
@@
Every point on the phase diagram of the SU(M)(M) Heisenberg model satisfies all EoMs 172, 165, 166 and 167 simultaneously. For the spin liquid phase we expect $u=0$ and $m=1$ for all $J$ and $\kappa$. On the other hand, the spin glass phase is characterized by $0<u\leq\kappa$ and $0<m<1$. The bound for $u$ arises from the constraint Eq. (76) and for integer $n$, $m$ is also an integer that describes the size of the diagonal blocks in the ansatz. In the limit $n\rightarrow 0$, $m$ becomes a continuous number between 0 and $1$ [5, 19]. In order to be able to find the correct solutions $\hat{g}(k)$ numerically for any point on the phase diagram, we need first to massage the EoMs into a more convenient form.
@@ B.9 Lorentzian two point function
The relation above is quite useful, as it will let us compute $g^{R}(\omega)$, from the values $\{u,m,g(k=0)\}$ that satisfy the triplet of Euclidean equations of motion 165, 166 and 167, derived above. In particular, the analytic continuation of (172) is the following:
@@ Appendix C Details on the spectral function
where the $\pm$-sign in the last line defines the Fermi-Dirac/Bose-Einstein distribution.
@@ Appendix D Analytic continuation of equations of motion
As we saw in 195, the retarded Green’s function is related to the euclidean Green’s function by an analytic continuation of the form $\omega\rightarrow-i\omega$. We can apply the same procedure to the Schwinger-Dyson equations to obtain the real time equations. which in turn allows us to find an expression for the self energy $\Sigma(\tau)$